OpenStudy (anonymous):

A taxicab moves five blocks due north, five blocks due east, and another two blocks due north. Assume all blocks are of equal size. What is the magnitude of the taxi’s displacement, start to finish?

OpenStudy (anonymous):



OpenStudy (anonymous):

lets see what we have here

OpenStudy (btaylor):

What is the magnitude of its displacement to the north (how many blocks did it travel N)? 7 blocks N What is the magnitude of its displacement to the east (how many blocks did it travel E)? 5 blocks E  use the pythagorean theorem: \[? = \sqrt{7^2 + 5^2} = \sqrt{49+25} = \sqrt{74} \approx 8.60 block\]